OpenStudy (anikate):

Write the expression using factorials. 45

OpenStudy (anikate):

help?

OpenStudy (anonymous):

\[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=(10!\div(8!\times2!))=45\]

OpenStudy (anikate):

uh... where did u get all that from MHM5000 does have to do anything with a formula

OpenStudy (anikate):

what's next the answer to that is -88.75

OpenStudy (anonymous):

I didn't get your Q! type it in regular English! I'm Starter in it! ;)

OpenStudy (anonymous):

The question is not clear.

OpenStudy (anikate):

write the expression using factorials. I need a factorial for 45 ex. 21*20*19=21P3

OpenStudy (anonymous):

Factorial for 45 as in something factorial = 45? It is not.

OpenStudy (anikate):

yup

OpenStudy (anonymous):

Factorial of a number = 1 * 2 * 3 * 4 * 5 *........* n

OpenStudy (anikate):

plz explain to me

OpenStudy (anonymous):

Nothing factorial = 45. But, ratios of some factorials can equal 45. There can be many.

OpenStudy (anikate):

ok

OpenStudy (anikate):

ok

OpenStudy (anonymous):

For example, 5! = 1*2*3*4*5 = 120 For example, 4! = 1*2*3*4 = 24 45 can't be represented as a factorial.

OpenStudy (anikate):

is there a decimal ?